6. Summarise the rules for writing of distribution of electrons in various shells for the first eighteen elements.

Grade:12th pass

1 Answers

Pawan Prajapati
3 years ago
1. The maximum number of electrons that can be accommodated in a shell is given by the formula : 2n2, where n = 1, 2, 3... 2. Maximum number of electrons in different shells are : K shell – n = 1; 2n2 = 2(1)2 = 2 L shell – n = 2; 2n2 = 2(2)2 = 8 M shell – n = 3; 2n2 = 2(3)2 = 18 N shell – n = 4; 2n2 = 2(4)2 = 32 3. Given number of electrons is 18. 4. The electrons are not taken unless the inner shells are filled. 5. Hence, the highest element has K – 2; L – 8; M – 8.